| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| rl: | 2024 081 14 44 31 Motor #3: appears stuck with a reference sensor activated. | 2 | 1 | B08224.194 | Toron | #194 | 05:26:33 | 1 | -1 | |
| rl: | 2024 081 14 45 43 New value set for watermark SUPPLY.VOLTAGE.HIGH[ | 8 | 3 | B08224.194 | Toron | #194 | 05:26:45 | 1 | -1 | |
| skc: | Azimuth drive turned OFF in op_st. | 1 | 2 | B08224.194 | Toron | #194 | 18:47:49 | 1 | -1 |
| Serial | YYYYJJJ | HG large | max correction |
|---|---|---|---|
| 194 | 2024082 | 32 | 4.9 |
| Serial | YYYYJJJ | HP large | max correction |
|---|---|---|---|
| 194 | 2024082 | 1 | 54.4 |
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| rl: | 2024 081 14 44 31 Checksum bad in low level command. | 1 | 3 | B08224.194 | Toron | #194 | 05:26:31 | 1 | -1 | |
| rl: | 2024 081 14 44 31 Motor #3: appears stuck with a reference sensor activated. | 1 | 3 | B08224.194 | Toron | #194 | 05:26:33 | 1 | -1 | |
| rl: | 2024 081 14 44 31 Motor #11: appears stuck with a reference sensor activated. | 1 | 3 | B08224.194 | Toron | #194 | 05:26:35 | 1 | -1 | |
| rl: | 2024 081 14 45 43 New value set for watermark SUPPLY.VOLTAGE.HIGH[ | 3 | 3 | B08224.194 | Toron | #194 | 05:26:45 | 1 | -1 | |
| rs: | HOME key pressed | 3 | 99 | B08224.194 | Toron | #194 | 12:05:43 | 1 | -1 | |
| pd: | Going to menu | 2 | 99 | B08224.194 | Toron | #194 | 12:06:12 | 1 | -1 | |
| re: | Zenith discrepancy = 37 | 1 | 99 | B08224.194 | Toron | #194 | 14:46:03 | 37 | -10 | 10 |
| dsp: | Doing dispersion test with Hg (UV port) | 2 | 99 | B08224.194 | Toron | #194 | 15:26:35 | 1 | -1 | |
| checkhg: | Last HG had to adjust the micrometer by at least 2 | 1 | 99 | B08224.194 | Toron | #194 | 17:37:47 | 1 | -1 |
| _jday | _loc | _serial | _skc | dt | ed | hg | hp | rs | sl |
|---|---|---|---|---|---|---|---|---|---|
| 082 | Toronto | #194 | Azimuth drive turned OFF in op_st.,menu,Must do HG now,Saving backup OP_ST.000,Starting schedule inside,Waiting until ZA = 170 now 127.1,Waiting until ZA = 170 now 128.3,Waiting until ZA = 170 now 129.3,Waiting until ZA = 170 now 130.3,Waiting until ZA = 170 now 131.2,Waiting until ZA = 170 now 132,Waiting until ZA = 170 now 132.8,Waiting until ZA = 170 now 133.4,Waiting until ZA = 170 now 133.9,Waiting until ZA = 170 now 134.3,Waiting until ZA = 170 now 134.7,Waiting until ZA = 170 now 134.9 | 5 | 2 | 39 | 33 | 4 | 5 |
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
|---|---|---|---|---|
| 194 | 31.9 | 6.6 |